[figure: 9.7 cm]

STRUCTURE OF SPUTTERED PERMALLOY/NOBLE METAL

MULTILAYERS WITH GMR EFFECT

B. Szymański Institute of Molecular Physics, Polish Academy of Sciences, Poznań, Poland

We report on structure of Py/NM multilayers, where Py = Ni_{83}Fe_{17} and NM = Cu, Ag and Au, obtained by face to face and magnetron sputtering onto various substrates. Although both Py and NM crystallize in fcc structures, they differ in the lattice constant and solubility. The multilayers showed GMR effect for appropriate thicknesses of the NM spacer. The structures of our films were investigated by the low and high angle Xray diffraction. The analysis of spectra revealed a high structural quality of multilayers. The samples were well textured polycrystalline films with very good periodicity in the direction perpendicular to the film plane and crystallites with a coherence length larger than the modulation wavelength. The differences observed in the low and high angle spectra for multilayers with various nonmagnetic spacer are discussed in relation to the GMR effect.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : B. Szymański

Address for correspondence : Institute of Molecular Physics, Polish Academy of Sciences PL 60-179 Poznań ul. Smoluchowskiego 17, Poland

Email address : szyman@ifmpan.poznan.pl